OpenStudy (anonymous):

What is the equation of the line that passes through (1, 2) and is parallel to the line whose equation is 4x + y + 1 = 0? 4x - y - 6 = 0 4x + y - 6 = 0 4x + y + 6 = 0

OpenStudy (welshfella):

first write the equation in the question in slope/intercept form which is y = -4x - 1 what is the slope of this line ?

OpenStudy (anonymous):

-4?

OpenStudy (welshfella):

yes now the line also goes through the point (1,2) so we put x1 = 1 and y1 = 2 in the equation y - y1 = -4(x - x1)

OpenStudy (welshfella):

y - 2 = -4(x - 1) rearrange this a nd you should get one of the options

OpenStudy (welshfella):

y - 2 = -4(x - 1) y - 2 = -4x + 4 can you continue?

OpenStudy (anonymous):

so 4x+y-6=0?

OpenStudy (welshfella):

yes